

********************************************************************
UARS WEEKLY STATUS REPORT - 4 Oct 2002
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all
available orbital tracking periods.
***********
SUSIM
***********
We are continuing to work with the UMPG Project, MPG, and FOT to
ensure that sufficient TDRSS telemetry time is available to allow
SUSIM to gather its daily scans under the new operational regime.
Otherwise, SUSIM successfully performed its daily mid-resolution and
MgII solar scans during this past week for periods of predicted UARS
telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in the UARS Monthly Status Review
meeting on October 3. At this meeting it was announced
that Juilo Marius is replacing Bill Guit as UARS
Mission Director.
The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 21 September 2002 (Orbit 60308, GMT Day 264)
through 27 September 2002 (Orbit 60413, GMT Day 270).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 265/0146 60324 POWER ON
HALOE 266/0256 60340 POWER OFF
HALOE 266/1314 60347 POWER ON
HALOE 267/1312 60364 POWER OFF
HALOE 269/2029 60396 POWER ON
The beta angle ranged between the angles of 56.1 degrees to 49.7
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 10 October 2002 (DOY 283).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.91 to 1.27 degrees C. Additional
battery performance data for each battery on 20 September and 27
September 2002 are provided in Appendix D.
The clock error ranged from -10.9 to 10.4 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
HALOE = 05
SOLSTICE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The hard disk on the prototype was temporarily fixed but is now
malfunctioning again. This problem should go away when the new hard
disk is installed this week. This will allow the FOT to test the
patches that have been made by the Eclipse development team. Once the
patches have been tested a new delivery of the software will be
installed. This will most likely not take place until the ERBS
integration is complete.
The FOT is also working hard in preparation for the college students
that will be in the control center to learn about spacecraft
operations. The FOT will try to formulate training materials and
structure a learning plan to teach the students. This is all part of
the new Educational Outreach initiative/contract, which starts on
October 1, 2002 for both UARS and TOMS.
MEETINGS
- None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 10 October 2002 (DOY 283).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 27 September 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
60322 264/2138 02-329 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
60324 265/0146 02-331 HALOE POWER ON
60340 266/0256 02-332 HALOE POWER OFF
60338 266/0004 02-333 SOLSTICE SOLS RESET CMDS
60347 266/1314 02-334 HALOE POWER ON
60355 267/0147 02-335 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
60364 267/1312 02-336 HALOE POWER OFF (TBAT TEMP)
60387 269/0615 02-337 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
60396 269/1948 02-338 HALOE POWER ON
60396 269/2029 02-338 HALOE POWER ON (COMPLETE)
60413 270/2323 02-339 OBC CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 27 September 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 27 September 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
60322 264/2140 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
60355 267/0150 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
60387 269/0616 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
60413 270/2326 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 27 September 2002
20 September 2002 - GMT Day 263 Beta = 54.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -140.0/-168.0 -3.05/-4.62 +00.0/+00.0 08.5 0.000 0.0
2 +11.2/-56.0 +4.51/+2.60 +22.4/-15.2 25.3 0.980 19.7
3 +16.8/-56.0 +3.23/+1.66 +22.0/-15.2 25.1 0.980 19.3
27 September 2002 - GMT Day 270 Beta = 49.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -145.6/-11.2 -3.68/-4.63 +00.0/+00.0 08.3 0.000 0.0
2 +11.2/-50.4 +5.14/+3.23 +24.4/-15.2 25.0 1.030 19.8
3 +22.4/-33.6 +3.87/+2.60 +24.0/-15.2 25.0 1.040 19.3
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 27 September 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-050 SOLSTICE 60336 265/2037 N/A SOLS UNEXPECTED STATE
02-051 FOT 60366 267/1951 30832 GENERIC LATE ACQUISITION
02-052 SOLSTICE 60366 267/1951 N/A LOADING PROBLEM
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-049 FOT 60293 263/0002 31477 NONSTANDARD BLIND ACQ
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-050 SOLSTICE 60336 265/2037 N/A SOLS UNEXPECTED STATE
02-051 FOT 60366 267/1951 30832 GENERIC LATE ACQUISITION
02-052 SOLSTICE 60366 267/1951 N/A LOADING PROBLEM